In the laboratory you dilute 5.27 mL of a concentrated 12.0 M nitric acid solution to a total volume of 50.0 mL. What is the concentration of the dilute solution?


______ M


1
Expert's answer
2021-10-17T11:24:51-0400

CM = n / V

n1 (HNO3) = n2 (HNO3)

n = CM x V

CM1 x V1 = CM2 x V2

12 x 0.00527 = CM2 x 0.0500

CM2 = (12.0 x 0.00527) / 0.0500 = 1.3 M